OpenStudy (anonymous):

Janine inflated 1 ball to a radius of 18 cm and another ball to a radius of 12 cm. How much more air was in the larger ball? Use 3.14 to approximate pi and express your answer in hundredths. cm^3

OpenStudy (anonymous):

@amistre64 @dmndlife24 @texaschic101 @Michele_Laino @JoannaBlackwelder @Thesmarterone @TheSmartOne

OpenStudy (anonymous):

volume of 18 cm ball - volume of 12 cm ball

OpenStudy (anonymous):

so pi18^2

OpenStudy (anonymous):

\[V=\frac{4}{3}*\pi*r^3\]

OpenStudy (anonymous):

1728x3.14x4/3

OpenStudy (anonymous):

5425.92 x 4/3

OpenStudy (anonymous):

can i turn 4/3 into a decimal?

OpenStudy (anonymous):

4/3 = 1 1/3 = 1.3333333333333333333333333 for infinity

OpenStudy (anonymous):

how u multiply it then?

OpenStudy (anonymous):

\[\frac{5425.92*4}{3}\]

OpenStudy (anonymous):

7234.56

OpenStudy (anonymous):

Right. Now we have to subtract this from the volume of the 18 cm ball.

OpenStudy (anonymous):

24416.64

OpenStudy (anonymous):

Right. So all we have left is to find the difference between them. In other words, 24416.64-7234.56=?

OpenStudy (anonymous):

17182.08

OpenStudy (anonymous):

And that'd be your answer: 17,182.08 cm^3

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

No problem :)